# Graph sub-sampling for divide-and-conquer algorithms in large networks

### v3

@@ 1 Introduction
Graphs, or networks [^1] We will use these terms interchangeably in this work., provide a simple and intuitive model for interconnected systems where objects or entities are represented as nodes and their relationships are denoted as edges. Networks have been applied to many real-world applications, ranging from friendships (Girvan and Newman,, 2002) to brain connectivity (Telesford et al.,, 2011). As their popularity has grown, so too has their size: it is not uncommon for researchers to now work with thousands, millions or even billions of nodes and edges (e.g., Backstrom et al.,, 2006; Rozemberczki et al.,, 2021). Community detection (Mukherjee et al.,, 2021; Fang et al.,, 2024; Cardoso et al.,, 2024; Chakrabarty et al.,, 2025), hypothesis testing (Ghoshdastidar and Von Luxburg,, 2018; Chakraborty et al.,, 2024) and link prediction (Zhang and Chen,, 2018; Kumar et al.,, 2020) are just a few examples where researchers must handle large graphs. Thus, modern network analysis methods must scale to such datasets if they are to be practically useful.
@@
Several recent works have proposed novel graph sampling algorithms for various tasks. Wang, (2023) develop an adaptive edge sub-sampling algorithm and apply it to find degree, clustering coefficient and K-core distributions. Similarly, Chen and Su, (2024) propose an edge-reinforced random walk sub-sampling strategy and show that sub-graphs extracted using the algorithm are similar to the original network, while Wang et al., (2024) propose a motif-based sub-sampler for supports estimation.
@@ 2.2 Divide-and-conquer for community structure
$\mathsf{PACE}$ begins by randomly sampling a sub-graph $S^{(b)}$ of graph $G$ with $qn$ nodes where $q\in(0,1)$ is the proportion of nodes in the sub-sample. With a sub-sample in hand, we can apply any community detection algorithm to this sub-sample, e.g., spectral (Ng et al.,, 2002), modularity maximization (Girvan and Newman,, 2002), semi-definite programming (Cai and Li,, 2015), etc. The resulting membership vectors of these clustering algorithms allow us to construct $\hat{\bf C}^{(b)}\in\{0,1\}^{n\times n}$ where $\hat{C}^{(b)}_{ij}=1$ if both nodes $i$ and $j$ appear in sub-graph $S^{(b)}$ and were assigned to the same community, and 0 otherwise. We repeat this process for $B$ sub-graphs, yielding $\hat{\bf C}^{(1)},\dots,\hat{\bf C}^{(B)}$.
@@
Compute ${\bf N}$ where $N_{ij}$ is the number of times nodes $i$ and $j$ were sampled together in a subgraph.  Compute $\hat{\bf C}$ where $\hat{C}_{ij}=\mathbb{I}(N_{ij}>\beta)\frac{\sum_{b=1}^{B}\hat{C}_{ij}^{(b)}}{N_{ij}}$  Perform $k$-means clustering on $\hat{\bf C}$ with $K$ clusters to obtain $\hat{\boldsymbol{c}}\in\{1,\dots,K\}^{n}$
@@
The main focus of this paper is the sub-sampling step of the divide-and-conquer algorithm, a topic that the authors of $\mathsf{PACE}$ give some attention too (see Section 2.3 of Mukherjee et al., (2021)). In particular, as possible sub-sampling schemes, they suggest: randomly sampling nodes; randomly sampling a node and including all of its neighbors within distance $h$ for some integer $h$ ($h$-hop neighborhoods); $1$-hop neighborhoods but removing the root node (ego neighborhoods); randomly sampling a node, including all of its neighbors, including all of the neighbors of the neighbors, etc. (onion neighborhoods); or sampling sub-graphs with roots at high degree nodes. For most experiments, the authors employ random node sampling.
@@ 2.3 Divide-and-conquer for core-periphery structure
We now turn our attention to the divide-and-conquer algorithm for detecting CP structures in large networks proposed by Yanchenko, (2022).
@@ 2.3.2 Divide-and-conquer algorithm
  Obtain $\hat{\boldsymbol{c}}$ as $\hat{c}_{i}=\sum_{b=1}^{B}\hat{c}_{i}^{(b)}/B$
@@ 2.4 Sub-graph sampling
In this method, nodes are again sampled at random, but now their probability of being sampled is proportional to their degree. Once the $qn$ nodes are sampled, then the sub-graph is completed with all associated edges between sampled nodes. Mathematically, this means sampling without replacement from a distribution with probability mass function $\mathsf{P}(Z=j)=d_{j}/\sum_{k=1}^{n}d_{k}$ where $d_{j}$ is the degree of node $j\in\{1,2,\dots,n\}$. As Leskovec and Faloutsos, (2006) point out, this approach will likely lead to sub-graphs which have a different degree distribution from the original graph, while also being denser.
@@
Breadth first search (BFS):
The previous three methods randomly sampled nodes or edges. For the remaining methods, we will consider graph exploration approaches. The BFS algorithm begins by sampling a node at random. Then all neighbors of this sampled node are included in the sub-graph. Next, all neighbors of the neighbors are also included in the sub-sample. This process continues until we have sampled $qn$ nodes.
Depth first search (DFS):
DFS can be considered as the complement of a BFS algorithm. This method also begins by randomly sampling a starting node. Then instead of including all neighbors of this starting node, it randomly chooses one neighbor. From there, it visits a neighboring node which has yet to be visited. This process continues until a node is reached such that all of its neighbors have already been visited. At this point, the algorithm goes back to the nearest unvisited node and continues the process again. The process continues until $qn$ nodes have been sampled. Krishnamurthy et al., (2005) studied the ability of BFS and DFS to preserve certain graph properties in the sub-samples, but found both performed poorly.
@@ 3.1 $\mathsf{PACE}$ algorithm
Corollary 3.1 (Corollary 3.1 from Mukherjee et al., (2021)): Assume that RN is used to construct the subgraphs. Then
@@ 3.2.1 General mis-classification result
where $\hat{\boldsymbol{c}}^{(S)},(\boldsymbol{c}^{*})^{(S)}\in\{0,1\}^{qn}$ are the estimated and true CP labels, respectively, restricted to sub-graph $S$ and $y_{i}^{(S)}=1$ if node $i$ is in the core and was not sampled in sub-graph $S$, and 0 otherwise.
@@ 3.2.2 Specific sub-sampling results
We assume networks are generated with a CP structure from the stochastic block model (SBM) (Holland et al.,, 1983). For adjacency matrix ${\bf A}$, let $A_{ij}\stackrel{{\scriptstyle ind.}}{{\sim}}\mathsf{Bernoulli}(\varrho_{n}P_{ij})$ where $P$ corresponds to a SBM that does not depend on $n$, and $\varrho_{n}=o(1)$ is a sparsity-inducing parameter. Furthermore, if $\boldsymbol{c}^{*}$ correspond to the true CP labels, then
$P_{ij}=\begin{cases}p_{11}&\text{if }c_{i}^{*}=c_{j}^{*}=1\\ p_{22}&\text{if }c_{i}^{*}=c_{j}^{*}=0\\ p_{12}&\text{otherwise}.\\ \end{cases}$
@@
$\lim_{n\to\infty}\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}y_{i}^{(S)}=\left(1-q_{n}\frac{p_{12}}{p_{22}}\right)\alpha_{n}.$ (8)
@@ 4 Simulation Study
In this section, we carry out extensive simulations of both divide-and-conquer algorithms. The goals of this section are to compare the performance of the algorithms under different sub-sampling schemes, as well as compare the divide-and-conquer algorithms against the original algorithm applied to the entire graph. Additionally, we hope to empirically validate the CP theoretical results. All experiments in this section were carried out in R on a Mac Mini with Apple M4 chip and 16 GB of memory parallelized using 9 cores. The code to replicate these simulations is available on the author’s GitHub: https://github.com/eyanchenko/subsamp.
@@ 4.1.1 Settings
$\gamma_{ij}(\hat{\boldsymbol{c}},\boldsymbol{c})=\begin{cases}1,&\hat{c}_{i}=\hat{c}_{j}\text{ and }c_{i}=c_{j}\\ 1,&\hat{c}_{i}\neq\hat{c}_{j}\text{ and }c_{i}\neq c_{j}\\ 0,&\text{otherwise}.\end{cases}$
@@
[…]
@@
1 $5000$ ${\bf(0.02,0.10)}$ $0.75$ $1000$ $250$
2 ${\bf(1000,8000)}$ $0.04$ $0.75$ $1000$ $250$
3 ${\bf(1000,8000)}$ $0.04$ $0.75$ ${\bf\frac{1}{2}}\boldsymbol{n}$ $250$
4 $5000$ $0.04$ ${\bf(0.50,0.95)}$ $1000$ $250$
[…]
6 $5000$ $0.05$ $0.75$ $1000$ ${\bf(100,500)}$
[…]
@@ 4.1.2 Results
The results are in Figures 1 and 2. In setting 1, save BFS and DFS, all methods yield an increasing ARI as $p_{11}$ increases. When $p_{11}\leq 0.04$, RNN has the largest ARI while RN does the best when $p_{11}>0.04$ among the divide-and-conquer methods. BFS improves at first before quickly dropping off, while DFS yields extremely low ARI values. Indeed, in all settings, it is clear that DFS performs poorly. Moreover, the algorithm applied to the full network yields the best performance when $p_{11}>0.03$. Turning to setting 2, the ARI increases with $n$ until around $n=4000$ before decreasing for most methods. By $n=8000$, these methods have experienced a drop off in performance. The full algorithm, on the other hand, has monotonically increasing performance with $n$, and yields the largest ARI values when $n\geq 4000$. When $B$ also increases with $n$ as in setting 3, DN, RE, RN, RNN and RW all have increasing ARI. RN and RNN perform slightly better and exceed or essentially match the ARI of Full for all $n$
The effects of the distribution of nodes between communities is evident in setting 4. Indeed, as the nodes are more disproportionately allocated between groups, each method decreases in performance with RN and Full being the most severely affected. As $B$ increases in setting 5, all methods, save DFS, see an improvement in performance. Interestingly, while RN and RNN plateau at an ARI of 1, DN, RE, RW and BFS all begin to yield slightly decreasing values of ARI for $B>2500$. Finally, setting 6 shows that as the number of nodes in the sub-samples increases, the communities are easier to detect for all methods except BFS and DFS. Among the divide-and-conquer methods, for most values of $q$, RN has the largest ARI followed by RNN. Note that the results for Full are constant in the last two settings since $B$ and $q$ do not appear as parameters in the full algorithm.
@@ 4.2.1 Settings
We again look at six simulation settings. For each setting, we generate a network, apply Algorithm 2 and compute the area under the receiver operator curve (AUC) between the outputted proportions and true labels. CP identification is fundamentally about classifying nodes as core or periphery and AUC is a popular metric for such tasks. The base CP detection algorithm is a greedy algorithm based on Algorithm 1 from Yanchenko, (2022). Similar to the community detection simulations, we also apply this algorithm to the entire network as a benchmark for comparison (Full). The networks are generated the same way as in the community detection simulations except now $p_{11}>p_{12}>p_{22}$ where $p_{11},p_{12},p_{22}$ represent the core-core, core-periphery, and periphery-periphery edge probabilities, respectively. For each setting, we fix $p_{22}=0.001$ and set $p_{12}=\frac{1}{2}p_{11}$.
@@
7 $5000$ ${\bf(0.002,0.020)}$ $0.01$ $1000$ $100$
8 ${\bf(1000,8000)}$ $0.004$ $0.01$ $1000$ $100$
9 ${\bf(1000,8000)}$ $0.004$ $0.01$ ${\bf\frac{1}{2}}\boldsymbol{n}$ $100$
10 $5000$ $0.004$ ${\bf(0.002,0.30)}$ $1000$ $100$
[…]
12 $5000$ $0.004$ $0.01$ $1000$ ${\bf(50,500)}$
[…]

### v4

@@ 1 Introduction
Graphs, or networks [^1] We will use these terms interchangeably in this work., provide a simple and intuitive model for interconnected systems where objects or entities are represented as nodes, and their relationships are denoted as edges. Networks have been applied to many real-world applications, ranging from friendships (Girvan and Newman,, 2002) to brain connectivity (Telesford et al.,, 2011). As their popularity has grown, so too has their size: it is not uncommon for researchers to now work with thousands, millions or even billions of nodes and edges (e.g., Backstrom et al.,, 2006; Rozemberczki et al.,, 2021). Community detection (Mukherjee et al.,, 2021; Fang et al.,, 2024; Cardoso et al.,, 2024; Chakrabarty et al.,, 2025), hypothesis testing (Ghoshdastidar and Von Luxburg,, 2018; Chakraborty et al.,, 2025) and link prediction (Zhang and Chen,, 2018; Kumar et al.,, 2020) are just a few examples where researchers must handle large graphs. Thus, modern network analysis methods must scale to such datasets if they are to be practically useful.
@@
Several recent works have proposed novel graph sampling algorithms for various tasks. Wang, (2023) develop an adaptive edge sub-sampling algorithm and apply it to find degree, clustering coefficient and $K$-core distributions. Similarly, Chen and Su, (2024) propose an edge-reinforced random walk sub-sampling strategy and show that sub-graphs extracted using the algorithm are similar to the original network, while Wang et al., (2024) propose a motif-based sub-sampler for support estimation.
@@ 2.2 Divide-and-conquer for community structure
$\mathsf{PACE}$ begins by randomly sampling a sub-graph $S^{(b)}$ of graph $G$ with $qn$ nodes where $q\in(0,1)$ is the proportion of nodes in the sub-sample. [^3] Note that $q$ may be chosen to ensure $qn$ is an integer, or $qn$ can simply be rounded to the nearest integer. With a sub-sample in hand, we can apply any community detection algorithm to this sub-sample, e.g., spectral (Ng et al.,, 2002), modularity maximization (Girvan and Newman,, 2002), semi-definite programming (Cai and Li,, 2015), etc. The resulting membership vectors of these clustering algorithms allow us to construct $\hat{\bf C}^{(b)}\in\{0,1\}^{n\times n}$ where $\hat{C}^{(b)}_{ij}=1$ if both nodes $i$ and $j$ appear in sub-graph $S^{(b)}$ and were assigned to the same community, and 0 otherwise. We repeat this process for $B$ sub-graphs, yielding $\hat{\bf C}^{(1)},\dots,\hat{\bf C}^{(B)}$.
@@
Compute ${\bf N}$ where $N_{ij}$ is the number of times nodes $i$ and $j$ were sampled together in a sub-graph.  Compute $\hat{\bf C}$ where $\hat{C}_{ij}=\mathbb{I}(N_{ij}>\beta)\displaystyle\frac{\sum_{b=1}^{B}\hat{C}_{ij}^{(b)}}{N_{ij}}$  Perform $k$-means clustering on $\hat{\bf C}$ with $K$ clusters to obtain $\hat{\boldsymbol{c}}\in\{1,\dots,K\}^{n}$
@@
The main focus of this paper is the sub-sampling step of the divide-and-conquer algorithm, a topic that the authors of $\mathsf{PACE}$ give some attention too (see Section 2.3 of Mukherjee et al., (2021)). In particular, as possible sub-sampling schemes, they suggest: randomly sampling nodes; randomly sampling a node and including all of its neighbors within distance $h$ for some integer $h$ ($h$-hop neighborhoods); $1$-hop neighborhoods but removing the root node (ego neighborhoods); randomly sampling a node, including all of its neighbors and all of the neighbors of the neighbors, etc. (onion neighborhoods); or sampling sub-graphs with roots at high degree nodes. For most experiments, the authors employ random node sampling.
@@ 2.3 Divide-and-conquer for core-periphery structure
We now turn our attention to the divide-and-conquer algorithm for detecting core-periphery (CP) structures in large networks proposed by Yanchenko, (2022).
@@ 2.3.2 Divide-and-conquer algorithm
Obtain $\hat{\boldsymbol{c}}$ as $\hat{c}_{i}=\displaystyle\frac{1}{B}\sum_{b=1}^{B}\hat{c}_{i}^{(b)}$
@@ 2.4 Sub-graph sampling
In this method, nodes are again sampled at random, but now their probability of being sampled is proportional to their degree. Once the $qn$ nodes are sampled, then the sub-graph is completed with all associated edges between sampled nodes. Mathematically, this means sampling without replacement from a distribution with probability mass function $\mathsf{P}(Z=j)=\displaystyle\frac{d_{j}}{\sum_{k=1}^{n}d_{k}}$ where $d_{j}$ is the degree of node $j\in\{1,2,\dots,n\}$. As Leskovec and Faloutsos, (2006) point out, this approach will likely lead to sub-graphs which have a different degree distribution from the original graph, while also being denser.
@@
Breadth-first search (BFS):
The previous three methods randomly sampled nodes or edges. For the remaining methods, we will consider graph exploration approaches. The BFS algorithm begins by sampling a node at random. Then all neighbors of this sampled node are included in the sub-graph. Next, all neighbors of the neighbors are also included in the sub-sample. This process continues until we have sampled $qn$ nodes. BFS algorithms have long been employed in the computer science and operations research communities, and can be carried out in linear time in terms of the number of nodes and edges (Cormen et al.,, 2022). The algorithm has been used for various tasks such as computing shortest paths (e.g., Bundy and Wallen,, 1984; Beamer et al.,, 2013; Burkhardt,, 2021), graph partitioning (Jovanovic et al.,, 2017) and finding connected components of a graph (Jain et al.,, 2017).
Depth-first search (DFS):
DFS can be considered as the complement of a BFS algorithm. This method also begins by randomly sampling a starting node. Then instead of including all neighbors of this starting node, it explores one (unexplored) edge to arrive at a new node. From here, a new unexplored edge is traversed, and this process continues until all of the current node’s edges have been visited. At this point, the algorithm returns to the nearest node with unexplored edges and the process restarts. This procedure continues until $qn$ nodes have been sampled. DFS also runs linearly in terms of the number of nodes and edges (Cormen et al.,, 2022) and has applications ranging from topological sorting of directed acyclic graphs and finding strongly connected components (Tarjan,, 1972), to multi-modal function optimization (Wang et al.,, 2022). Finally, Krishnamurthy et al., (2005) studied the ability of BFS and DFS to preserve certain graph properties in the sub-samples, but found both performed poorly.
@@
Please see the Supplemental Materials for a figure which compares each of these algorithms on a small graph.
@@ 3.1 $\mathsf{PACE}$ algorithm
Corollary 3.1 (Corollary 3.1 from Mukherjee et al., (2021)): Assume that RN is used to construct the sub-graphs. Then
@@ 3.2.2 Specific sub-sampling results
We assume networks are generated with a CP structure from the stochastic block model (SBM) (Holland et al.,, 1983). For an adjacency matrix ${\bf A}$, let $A_{ij}\stackrel{{\scriptstyle ind.}}{{\sim}}\mathsf{Bernoulli}(\varrho_{n}P_{ij})$ where $P$ corresponds to a SBM that does not depend on $n$, and $\varrho_{n}=o(1)$ is a sparsity-inducing parameter. Furthermore, if $\boldsymbol{c}^{*}$ corresponds to the true CP labels, then
$P_{ij}=\begin{cases}p_{11}&\text{if }c_{i}^{*}=c_{j}^{*}=1,\\ p_{22}&\text{if }c_{i}^{*}=c_{j}^{*}=0,\\ p_{12}&\text{otherwise}.\\ \end{cases}$
@@
$\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}y_{i}^{(S)}\sim\left(1-q_{n}\frac{p_{12}}{p_{22}}\right)\alpha_{n},$ (8)
where $\sim$ denotes the asymptotic behavior for large $n$.
@@ 4 Simulation Study
In this section, we carry out extensive simulations of both divide-and-conquer algorithms. The goals of this section are to compare the performance of the algorithms under different sub-sampling schemes, as well as compare the divide-and-conquer algorithms against the original algorithm applied to the entire graph. Additionally, we hope to empirically validate the CP theoretical results. All experiments in this section were carried out in R (Version 4.4.2) on a Mac Mini with Apple M4 chip and 16 GB of memory parallelized using 9 cores. The code to replicate these simulations is available on the author’s GitHub: https://github.com/eyanchenko/subsamp.
@@ 4.1.1 Settings
$\gamma_{ij}(\hat{\boldsymbol{c}},\boldsymbol{c})=\begin{cases}1&\hat{c}_{i}=\hat{c}_{j}\text{ and }c_{i}=c_{j},\\ 1&\hat{c}_{i}\neq\hat{c}_{j}\text{ and }c_{i}\neq c_{j},\\ 0&\text{otherwise}.\end{cases}$
@@
An overview of the simulation settings is available in Table 1. Setting 1 looks at the effect of the intra-community edge probability where larger values indicated stronger community structure, and we vary $p_{11}=0.020$, $0.025$, $\dots$, $0.10$. Settings 2 and 3 increase the number of nodes in the network with $B$ held fixed and increasing with $n$, respectively, for $n=1\hskip 1.9919pt000$, $2\hskip 1.9919pt000$, $\dots$, $8\hskip 1.9919pt000$. The proportion of nodes, $\alpha$, in the first community is the subject of setting 4 where greater imbalance in the community sizes should make the task more difficult. We consider $\alpha=0.50$, $0.60$, $\dots$, $0.90$, $0.95$. Finally, settings 5 and 6 look at the hyper-parameters in the $\mathsf{PACE}$ model, namely the number of sub-graphs drawn and size of sub-graphs, respectively. In setting 7, $B$ varies from $B=100$, $250$, $500$, $1\hskip 1.9919pt000$, $2\hskip 1.9919pt500$, $5\hskip 1.9919pt000$, $10\hskip 1.9919pt000$; while in setting 8, $qn$ varies from $qn=100$, $200$, $\dots$, $500$. For all parameter combinations, we generate 100 networks and record the average ARI and total computation time. In the Supplemental Materials, we also record the average computation time to detect the community structure on each sub-sample. Finally, we use the Fast Greedy algorithm (Clauset et al.,, 2004) as the base algorithm as in Zhang et al., (2023). We also apply this algorithm to the entire network as a competing method to the divide-and-conquer approaches, labeled Full. In the Supplemental Materials, we repeat the same experiments using Walktrap (Pons and Latapy,, 2005) as the base algorithm, while also reporting the average sub-graph density for each algorithm across all settings.
@@
1 $5\hskip 1.9919pt000$ ${\bf(0.02,0.10)}$ $0.75$ $1\hskip 1.9919pt000$ $250$
2 ${\bf(1\hskip 1.9919pt000,8\hskip 1.9919pt000)}$ $0.04$ $0.75$ $1\hskip 1.9919pt000$ $250$
3 ${\bf(1\hskip 1.9919pt000,8\hskip 1.9919pt000)}$ $0.04$ $0.75$ ${\bf\frac{1}{2}}\boldsymbol{n}$ $250$
4 $5\hskip 1.9919pt000$ $0.04$ ${\bf(0.50,0.95)}$ $1\hskip 1.9919pt000$ $250$
5 $5\hskip 1.9919pt000$ $0.04$ $0.75$ ${\bf(100,10\hskip 1.9919pt000)}$ $250$
6 $5\hskip 1.9919pt000$ $0.05$ $0.75$ $1\hskip 1.9919pt000$ ${\bf(100,500)}$
Table 1: Simulation settings for $\mathsf{PACE}$ divide-and-conquer algorithm. $n$ is the number of nodes in the network and for settings 2 and 3 it varies from $n=1\hskip 1.9919pt000$, $2\hskip 1.9919pt000$, $\dots$, $8\hskip 1.9919pt000$; $p_{11}$ is the intra-community edge probability and for setting 1 it varies from $p_{11}=0.020$, $0.025$, $\dots$, $0.10$; $\alpha$ is the proportion of nodes in community 1 and in setting 4 varies from $\alpha=0.50$, $0.60$, $\dots$, $0.90$, $0.95$; $B$ is the number of sub-graphs sampled and in setting 7 varies from $B=100$, $250$, $500$, $1\hskip 1.9919pt000$, $2\hskip 1.9919pt500$, $5\hskip 1.9919pt000$, $10\hskip 1.9919pt000$; and $qn$ is the number of nodes per sub-graph and in setting 8 varies from $qn=100$, $200$, $\dots$, $500$. The parameters which vary are highlighted in ${\bf bold}$.
@@ 4.1.2 Results
The results are in Figures 1 and 2. In setting 1, save BFS and DFS, all methods yield an increasing ARI as $p_{11}$ increases. When $p_{11}\leq 0.04$, RNN has the largest ARI while RN does the best when $p_{11}>0.04$ among the divide-and-conquer methods. BFS improves at first before quickly dropping off, while DFS yields extremely low ARI values. Indeed, in all settings, it is clear that DFS performs poorly. Moreover, the algorithm applied to the full network yields the best performance when $p_{11}>0.03$. Turning to setting 2, the ARI increases with $n$ until around $n=4\hskip 1.9919pt000$ before decreasing for most methods. By $n=8\hskip 1.9919pt000$, these methods have experienced a drop off in performance. The full algorithm, on the other hand, has monotonically increasing performance with $n$, and yields the largest ARI values when $n\geq 4\hskip 1.9919pt000$. When $B$ also increases with $n$ as in setting 3, DN, RE, RN, RNN and RW all have increasing ARI. RN and RNN perform slightly better and exceed or essentially match the ARI of Full for all $n$
The effects of the distribution of nodes between communities is evident in setting 4. Indeed, as the nodes are more disproportionately allocated between groups, each method decreases in performance with RN and Full being the most severely affected. As $B$ increases in setting 5, all methods, save DFS, see an improvement in performance. Interestingly, while RN and RNN plateau at an ARI of 1, DN, RE, RW and BFS all begin to yield slightly decreasing values of ARI for $B>2\hskip 1.9919pt500$. Finally, setting 6 shows that as the number of nodes in the sub-samples increases, the communities are easier to detect for all methods except BFS and DFS. Among the divide-and-conquer methods, for most values of $q$, RN has the largest ARI followed by RNN. Note that the results for Full are constant in the last two settings since $B$ and $q$ do not appear as parameters in the full algorithm.
@@ 4.2.1 Settings
We again look at six simulation settings. For each setting, we generate a network, apply Algorithm 2 and compute the area under the receiver operator curve (AUC) between the outputted proportions and true labels. CP identification is fundamentally about classifying nodes as core or periphery and AUC is a popular metric for such tasks (e.g., Kojaku and Masuda,, 2018; Naik et al.,, 2021). The base CP detection algorithm is a greedy algorithm based on Algorithm 1 from Yanchenko, (2022). Similar to the community detection simulations, we also apply this algorithm to the entire network as a benchmark for comparison (Full). The networks are generated the same way as in the community detection simulations except now $p_{11}>p_{12}>p_{22}$ where $p_{11},p_{12},p_{22}$ represent the core-core, core-periphery, and periphery-periphery edge probabilities, respectively. For each setting, we fix $p_{22}=0.001$ and set $p_{12}=\frac{1}{2}p_{11}$.
@@
7 $5\hskip 1.9919pt000$ ${\bf(0.002,0.020)}$ $0.01$ $1\hskip 1.9919pt000$ $100$
8 ${\bf(1\hskip 1.9919pt000,8\hskip 1.9919pt000)}$ $0.004$ $0.01$ $1\hskip 1.9919pt000$ $100$
9 ${\bf(1\hskip 1.9919pt000,8\hskip 1.9919pt000)}$ $0.004$ $0.01$ ${\bf\frac{1}{2}}\boldsymbol{n}$ $100$
10 $5\hskip 1.9919pt000$ $0.004$ ${\bf(0.002,0.30)}$ $1\hskip 1.9919pt000$ $100$
11 $5\hskip 1.9919pt000$ $0.004$ $0.01$ ${\bf(100,10\hskip 1.9919pt000)}$ $100$
12 $5\hskip 1.9919pt000$ $0.004$ $0.01$ $1\hskip 1.9919pt000$ ${\bf(50,500)}$
Table 2: Simulation settings for CP divide-and-conquer algorithm. $n$ is the number of nodes in the network and for settings 2 and 3 it varies from $n=1\hskip 1.9919pt000$, $2\hskip 1.9919pt000$, $\dots$, $8\hskip 1.9919pt000$; $p_{11}$ is the core-core edge probability and for setting 1 it varies from $p_{11}=0.002$, $0.003$, $\dots$, $0.020$; $\alpha$ is the proportion of nodes in the core and in setting 4 varies from $\alpha=0.002$, $0.004$, $\dots$, $0.008$, $0.01$, $0.02$, $\dots$, $0.05$, $0.10$, $0.20$, $0.30$; $B$ is the number of sub-graphs sampled and in setting 7 varies from $B=100$, $250$, $500$, $1\hskip 1.9919pt000$, $2\hskip 1.9919pt500$, $5\hskip 1.9919pt000$, $10\hskip 1.9919pt000$; and $qn$ is the number of nodes per sub-graph and in setting 8 varies from $qn=50$, $100$, $250$, $500$. The parameters which vary are highlighted in ${\bf bold}$.
The simulation settings are summarized in Table 2. In setting 7, we vary $p_{11}=0.002$, $0.003$, $\dots$, $0.020$ where larger values correspond to a stronger CP structure. Settings 8 and 9 look at the effect of increasing the network size where the former fixes $B$ while the latter allows $B$ to increase with $n$, i.e., $B=\frac{1}{2}n$. Both settings consider $n=1\hskip 1.9919pt000$, $2\hskip 1.9919pt000$, $\dots$, $8\hskip 1.9919pt000$. The effect of the core size, $\alpha$, is studied in setting 10 for $\alpha=0.002$, $0.004$, $\dots$, $0.008$, $0.01$, $0.02$, $\dots$, $0.05$, $0.10$, $0.20$, $0.30$. The choice of hyper-parameters is investigated in settings 11 and 12, varying the values of $B$ and $q$, respectively. We vary $B=100$, $250$, $500$, $1\hskip 1.9919pt000$, $2\hskip 1.9919pt500$, $5\hskip 1.9919pt000$, $10\hskip 1.9919pt000$ and $qn=50$, $100$, $250$, $500$ in these settings, respectively. Each setting is repeated for 100 Monte Carlo replications and the average AUC and total run time are reported. In the Supplemental Materials, we also report the average detection time per sub-sample.